FARO Technologies, whose products include portable 3D measurement equipment, recently had 17 million shares outstanding trading at $35 a share. Suppose the company announces its intention to raise $200 million by selling new shares.

a. How large a gain or loss in aggregate dollar terms do market signaling studies suggest existing FARO shareholders will experience on the announcement date? Can you specify how you got the percentage for the expected loss, please?

b. What percentage of the value of FARO’s existing equity prior to the announcement is this expected gain or loss?

d. At what price should FARO expect its existing shares to sell immediately after the announcement?
Answer #1

a) It is expected that there would be loss for each shareholder considering that the company lists the new shares at the market price of $35. Assuming that there is no expansion in the business and revenue remaining the same, the net income will now be distributed to greater number of shareholder hence the price will decrease.

Considering $200 M of capital raise at $35 each, the company will issue $(200/35)=5.71 M more shares. The market capitalization of $35 * 17 M =$595 M will be now divided to 17M + 6M = 23 M shares that is price per share is $595 M/23 M = $25.87.

Expected loss percentage= $35-$25.87/$35*100=26.10%

Expected loss=26%*$200M=$52M

b) Answer= $52M/$585M*100=8.9%

d) The price would drop to $25.87 per share after the announcement.